Dr. M. B. Goshe PROBLEM SET 6 BCH 35 4. (0.5 pts) During the last step in the preparation of Bearnaise sauce, melted butter is incorporated into a liquid, which contains wine, vinegar, and egg yolks. The egg yolks are needed to stabilize the sauce and avoid separation; the stabilizing agent in the egg yolks is lecithin (Phosphatidylcholine),. Suggest why the egg yolks are necessary 5. (0.5 pts) The carbon-carbon bond distance for single-bonded carbons such as those in a saturated fatty acyl chain is about 1 its fully extended form. If two molecules of palmitate were placed end to end, how would their total lengt thickness of a lipid bilayer is 50 to 80 Å.) .5 A. Estimate the length of a single molecule of palmitate in h compare with the thickness of the lipid bilayer in a biological membrane? (The


help with these please!
Answer #1

Ans. #1. Phosphatidylcholine is a type of phospholipid, being amphipathic in nature.

The nature of molecules with hydrophobic region to form stable assemblies/aggregates in an aqueous medium is called ‘hydrophobic effect’. Hydrophobic molecules in butter (liquid) can’t form stable aggregates in water on their own. They disperse over aqueous layer and the form thin film when allowed to stand for a while.

The

The hydrophobic tail of several phosphatidylcholine interacts with hydrophobic molecules of butter whereas their polar glycerol-phosphate head simultaneously interacts with polar aqueous environment at the periphery forming micelles. The micelles are stable entity and don’t let the butter come out of solution when allowed to rest.

Therefore egg yolk is required to stabilize butter in the sauce (like natural stabilization of milk fat in milk).

#2. Palmitic acid = CH₃(CH₂)₁₄COOH

Number of carbon atom = 16

No. of C-C bonds = total number of C-atoms – 1 = 16 – 1 = 15

C-C bond length = 1.5A

Total length of molecule when fully extended = Number of C-C bonds x C-C bond length

                                                = (1.5 A) x 15

                                                = 22.5 A

Note: the length of C-O bond is ignored as no information is provided regarding this.

# Total length of two palmitate molecules when aligned end to end =

                                                2 x extended length of one palmitate molecule

                                                2 x 22.5 A

                                                = 45 A

Comparison: It is in close approximation of the thickness of lipid bilayer. The lipid bilayer consists of the hydrophobic core (two palmitate molecules lying end to end) and the polar heads (glycerol-phosphate) on each face. When the dimension of polar head is accounted, the total thickness of lipid bilayer would fall in the given range.